OpenStudy (anonymous):

a couple have 5 children, what is the probability that they are all girls?

OpenStudy (anonymous):

\[(1/2)^5 = 1/32\]

OpenStudy (anonymous):

1/32 is also the probability of having all boys. Or having any combination of boys and girls. This is because there are 32 possible combinations for boys and girls, each with the same probability.

OpenStudy (lgbasallote):

unless there is a third gender, there are two main genders so the probability first child is girl is 1/2 probability second child is girl too 1/2 x 1/2 third child is girl 1/2 x 1/2 x 1/2 fourth child 1/2 x 1/2 x 1/2 x 1/2 fifth child 1/2 x 1/2 x 1/2 x 1/2 x 1/2 got it? :D

OpenStudy (anonymous):

Thank you so much! I finally understand these questions!

OpenStudy (anonymous):

wait..do you have the answer towards it? I don't that it's that straight forward..

OpenStudy (anonymous):

at the back of the book it gives the answers and it says 1/32 so it was correct the way the first person did it :D

OpenStudy (lgbasallote):

lol the second person's answer is also 1/32 too ;)

OpenStudy (lgbasallote):

1/2 x 1/2 x 1/2 x 1/2 x 1/2 = 1/32

OpenStudy (anonymous):

really? sorry Igbasallote i just seen the first one :) im new to this i just signed up litrally 2 secs ago so i dunno anything..how does the website work?

OpenStudy (lgbasallote):

when you're confused in something then ask a question. then we teach you how to get the answer and solution. then when you see other questions you know, teach as well! :)

OpenStudy (anonymous):

I love this website already! my books are so confusing no proper examples in them what-so-ever, thanks anyway for explaining how to do the question..really helped! :D

OpenStudy (lgbasallote):

\[\huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]